To what height should a rectangular cylinder having square base length 10cm be filled so that the total force on the bottom is equal to the one of the side

Grade:11

1 Answers

Arun
6 years ago
Side pressure can be assumed at the COM of side wall so for equal force let the side of square base is aBottom force = pgh(a)²Side wall pressure = (pgh/2)(4ah)According given condition both are equal So h = a/2 and it is equal to half of side of rectangular base